OpenStudy (anonymous):

ln 3x + ln 2x = 9 Solve for x

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \ln(3x) + \ln(2x) = 9 }\) \(\large\color{black}{ \displaystyle \ln(3x\times2x) = 9 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \ln(6x^2) = 9 }\) \(\large\color{black}{ \displaystyle \ln(6x^2) = 9\ln(e) }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \ln(6x^2) = \ln(e^9) }\) \(\large\color{black}{ \displaystyle 6x^2=e^9 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x^2=\frac{e^9}{6} }\)

OpenStudy (solomonzelman):

and you can do the rest

OpenStudy (anonymous):

Thanks man!